a round target has 3 circles with radii of 2cm, 10cm, & 20cm. what is the probability that a dart randomly thrown hits the target within the outer ring? (outer ring's radii is 20cm)
Answer by checkley77(12569)

A(20)=pir^2
A=3.14*20^2
A=3.14*400
A=1,256 total area
A(10)=3.14*10^2
A=3.14*100
A=314 area inside the 10 circle.
(1,256-314)/1,256=942/1,256=.75 or 75% chance of hitting the 20 ring.